OpenStudy (anonymous):

Can someone please help me? I don't understand Stoichiometry, and I need someone to explain this to me and how to do it... How many grams of aluminum chloride must decompose to produce 78.3 milliliters of aluminum metal, if the density of aluminum is 2.70 g/mL? Show all steps of your calculation as well as the final answer. AlCl3 → Al + Cl2

OpenStudy (nechirwan):

you first be balance the equation 2AlCl3-->2Al+3Cl2 as we see the mole of Al is equal to AlCl3 =2 there fore by calculating the mole of Al we know mole of AlCl3 density=weight/volume weight =density x volume =2.70 g/mL x78.3 ml =211.41 gm mole =wt/m.wt=211.41/27=7.83 mole Al =AlCl3 mole AlCl3= wt/m.wt=>wt=m.wt x mole =98 x7.83=767.34 gm

OpenStudy (anonymous):

I don't understand any of that... ;~;

OpenStudy (nechirwan):

it is easy just you must balance equation

OpenStudy (anonymous):

How do you balance an equation to begin with?

OpenStudy (nechirwan):

AlCl3 → Al + Cl2 in this equation we have 3 chlorine in the left but we have 2 chlorine in the right there fore U must multiply left by 2 and right by 3

OpenStudy (anonymous):

Then I'd have 6(?)

OpenStudy (nechirwan):

yes in both side

OpenStudy (anonymous):

I really don't understand...

OpenStudy (nechirwan):

you must study and solve more problem

OpenStudy (anonymous):

.... Okay?

OpenStudy (aaronq):

www.khanacademy.org